OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region bounded by y=x^3 and x=y^3 in the first quadrant about the x-axis.

terenzreignz (terenzreignz):

Can you set up the definite integral that leads to this answer?

OpenStudy (anonymous):

no im totally lost

OpenStudy (anonymous):

it is multiple choice

terenzreignz (terenzreignz):

There must be limits of sorts...

terenzreignz (terenzreignz):

from 0 to 1, perhaps?

OpenStudy (anonymous):

that is the only thing the question says

terenzreignz (terenzreignz):

I see. Well, I'll assume it's from 0 to 1, then, because that is where they intersect, ok? :)

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

Well, you realise that from 0 to 1, x = y^3 is higher up than y = x^3, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the answer can be 16/35pi, 16/7pi, 18/35, 7/2, or 16pi

terenzreignz (terenzreignz):

In other words, we write both of them as functions of x... like so \[f(x) = \sqrt[3]{x}\] and \[g(x) = x^{3}\] And from 0 to 1, f(x) is greater than g(x)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what do we do next?

terenzreignz (terenzreignz):

If f(x) is always greater than or equal to g(x) in the interval, then the volume of the solid of rotation is given by this formula (I can do my best to derive it if you so wish) \[\pi \int\limits_{a}^{b}[f(x)^{2}-g(x)^{2}]dx\]

terenzreignz (terenzreignz):

Where (a,b) is your interval :) Can you evaluate the integral?

OpenStudy (anonymous):

yes please

terenzreignz (terenzreignz):

yes please what?

OpenStudy (anonymous):

oh that was it.. what do we do from there?:

terenzreignz (terenzreignz):

Well, what's the interval, it's [0,1], right? and we have our values for f and g so all that's left is to plug them in :)

OpenStudy (anonymous):

how do we plug a and b into x?

terenzreignz (terenzreignz):

your values for a and b are 0 and 1, respectively, as we have that interval :)

terenzreignz (terenzreignz):

Are you evaluating it?

OpenStudy (anonymous):

\[\int \left(\pi \sqrt[3]{x}^2-\pi \left(x^3\right)^2\right) \, dx=\pi \left(\frac{3 x^{5/3}}{5}-\frac{x^7}{7}\right)+K \]

OpenStudy (anonymous):

Evaluate the above from 0 to 1 to get \[ \int_0^1 \left(\pi \sqrt[3]{x}^2-\pi \left(x^3\right)^2\right) \, dx=\frac{16 \pi }{35} \]